# Selective “Selective Prediction”: Reducing Unnecessary Abstention in Vision-Language Reasoning

### v1

@@ 5.1 Reliability and Relevance Ablations
$\mathcal{M}_{\texttt{QGen}}$ $\mathcal{R}$($\downarrow$) $\Phi_{1}$($\uparrow$) $\mathcal{C}$($\uparrow$) $\mathbb{R}_{\texttt{SP}}$($\uparrow$)
@@
Image Initial prediction {LLM Question $q_{j}$} {VLM Answer $a_{j}$} $\pi_{\texttt{{VLM}}}(a_{j})$]
@@
We perform an ablation of the reliability and relevance components of ReCoVERR   for calibrated BLIP2 and InstructBLIP, at a risk tolerance of 20%. In the ReCoVERR formulation, for an evidence $e_{j}$ to be considered reliable, the VLM confidence $\pi_{\texttt{{VLM}}}(a_{j}|I,Q_{j})$ must be at least 1r1-r, and for it to be considered relevant, the defeasible relevance $\delta(e_{j})$ must be at least $\delta_{\texttt{min}}=0.2$. For the reliability ablation, we lower the minimum evidence confidence to be $0.5$ instead of 1r=0.81-r=0.8. For the relevance ablation, we set $\delta_{\texttt{min}}=0$.
Table 2 shows the results of the reliability and relevance ablation. We see that loosening the evidence confidence bound to be $0.5$ instead of 1r1-r causes a marked increase in risk for both VLMs, with system risk crossing the risk tolerance significantly. Ablating the relevance component results in a decrease in recall and effective reliability for InstructBLIP, but virtually no effect for BLIP2.
@@ 6 Conclusion and Related Works
A large breadth of prior work has studied the ability of models to abstain from answering Chow (1957); De Stefano et al. (2000); El-Yaniv et al. (2010). This ability is especially important when the model’s prediction cannot be trusted Jiang et al. (2018), particularly for OOD Kamath et al. (2020) and adversarial Varshney et al. (2022) inputs.
In the multimodal domain, previous works have explored abstention for error recovery Wu et al. (2023); Khan and Fu (2023). Whitehead et al. (2022) establish the ReliableVQA framework, wherein VQA models have the option to abstain from answering. Shukor et al. (2023) evaluates VQA models on their ability to abstain when a question does not apply to the image. Dancette et al. (2023) trains VQA models to abstain for out-of-distribution inputs. In contrast to learning to abstain, we reduce abstention of multimodal selective prediction while maintaining reliability.
@@ Appendix A Self-Prompting for Producing Calibrated VLM Confidence Estimates
A.1 Comparison Against Other Confidence Measures
Figure 6: Calibration curves and expected calibration error (ECE) for various VLM confidence measures, using both BLIP2 and InstructBLIP, on A-OKVQA questions.
We compare our Self-Prompting approach for extracting VLM confidences to several other confidence measures that look at token probabilities of the VLM’s prediction. Suppose, for a question $Q$ about an image $I$, the VLM generates an answer $a=(a_{1},a_{2},...,a_{N})$ consisting of $N$ tokens. We consider the following confidence measures: Product of Token Probabilities $\Big(\prod_{i=1}^{N}P_{\texttt{VLM}}(a_{i}|I,Q,a_{1...i-1})\Big)$, Mean of Token Probabilities $\Big(\frac{1}{N}\sum_{i=1}^{N}P_{\texttt{VLM}}(a_{i}|I,Q,a_{1...i-1})\Big)$ and First Token Probability $\Big(P_{\texttt{VLM}}(a_{1}|I,Q)\Big)$. Figure 6 shows calibration curves for the above three confidence measures, as well as Self-Prompting. We find that Self-Prompting has lower calibration error than the other three confidence measures. Further, looking at the confidence distributions, we see that the other three confidence measures result in the majority of predictions having very low confidence (<0.5<0.5), and these under-confident predictions are actually more accurate than their confidence indicates.
A.2 Platt Scaling
@@ Appendix B Prompts for Text-Only Language Models
Target question: $\{Q\}$. Generate $K$ sub-questions that might help you confirm whether the answer to the target question is $\{a\}$’.
@@
Checking evidence sufficiency $\mathcal{M}_{\texttt{NLI}}$ $\mathcal{H}$: Hypothesis statement
$\mathcal{E}_{RR}$: A list of reliable and relevant evidences $e_{1},e_{2},$...]
Premise: $\{\text{evidence statements in }\mathcal{E}_{RR}\text{ concatenated together}\}$
@@ Appendix C OK-VQA and Sherlock Task Details
Is the given inference true for the given image? {Inference} Options: yes, no
We observe poor calibration on these tasks, possibly due to the out-of-distribution nature of images with region boxes drawn on, due to which we do not perform ReCoVERR experiments by calibrating on these respective datasets.
Risk tolerance = 10% Risk tolerance = 20%
VLM ReCoVERR Risk ReCoVERR Coverage Vanilla SelPred Coverage @ ReCoVERR Risk ReCoVERR Risk ReCoVERR Coverage Vanilla SelPred Coverage @ ReCoVERR Risk
BLIP2 (Off-the-Shelf) 14.3 26.0 21.0 21.7 47.3 44.9
BLIP2 (Calibrated) 14.0 23.6 23.3 16.1 36.7 29.2
InstructBLIP (Off-the-Shelf) 10.9 33.5 33.3 17.9 63.7 58.5
InstructBLIP (Calibrated) 11.8 38.7 32.3 18.6 65.0 59.5
Table 7: Coverage of ReCoVERR versus vanilla selective prediction, at the final resultant risk that ReCoVERR produces. Measurements in blue indicate ReCoVERR’s coverage exceeding that of vanilla selective prediction, at the same risk.
Appendix D Comparing ReCoVERR and Vanilla Selective Prediction at ReCoVERR’s Risk
We observe in Table 1 that ReCoVERR has both higher coverage and higher risk than vanilla selective prediction. However, one simple way to increase the vanilla selective prediction coverage is to set the threshold $\gamma_{@r}$ at a risk level that corresponds to ReCoVERR’s risk. Table 7 shows the coverage of vanilla selective prediction and ReCoVERR at the same risk level as ReCoVERR’s risk.

### v2

@@ 5.1 Reliability and Relevance Ablations
We perform an ablation of the reliability and relevance components of ReCoVERR   for calibrated BLIP2 and InstructBLIP, at a risk tolerance of 20%. In the ReCoVERR formulation, for an evidence $e_{j}$ to be considered reliable, the VLM confidence $\pi_{\texttt{{VLM}}}(a_{j}|I,Q_{j})$ must be at least $1-r$, and for it to be considered relevant, the defeasible relevance $\delta(e_{j})$ must be at least $\delta_{\texttt{min}}=0.2$. For the reliability ablation, we lower the minimum evidence confidence to be $0.5$ instead of $0.81-r=$ 0.8. For the relevance ablation, we set $\delta_{\texttt{min}}=0$.
@@
[…]
@@
Image Initial prediction {LLM Question $q_{j}$} {VLM Answer $a_{j}$} [$\pi_{\texttt{{VLM}}}(a_{j})$]
@@
Table 3 shows the results of the reliability and relevance ablation. We see that loosening the evidence confidence bound to be $0.5$ instead of $1-r$ causes a marked increase in risk for both VLMs, with system risk crossing the risk tolerance significantly. Ablating the relevance component results in a decrease in recall and effective reliability for InstructBLIP, but virtually no effect for BLIP2.
@@ 6 Conclusion and Related Works
A large breadth of prior work has studied the ability of models to abstain from answering Chow (1957); De Stefano et al. (2000); El-Yaniv et al. (2010). This ability is especially important when the model’s prediction cannot be trusted Jiang et al. (2018), particularly for OOD Kamath et al. (2020) and adversarial Varshney et al. (2022) inputs. Similar to our work, low coverage caused by low risk tolerance has also been observed and mitigated in text-only selective prediction systems Varshney and Baral (2023).
In the multimodal domain, previous works have explored abstention for error recovery Wu et al. (2023); Khan and Fu (2023). Whitehead et al. (2022b) establish the ReliableVQA framework, wherein VQA models have the option to abstain from answering. Shukor et al. (2023) evaluates VQA models on their ability to abstain when a question does not apply to the image. Dancette et al. (2023) trains VQA models to abstain for out-of-distribution inputs. In contrast to learning to abstain, we reduce abstention of multimodal selective prediction while maintaining reliability.
@@
7 Acknowledgements
Jack Hessel partially completed this work while at AI2.
@@ Appendix A Self-Prompting for Producing Calibrated VLM Confidence Estimates
A.1 Platt Scaling
@@ Appendix B Prompts for Text-Only Language Models
Target question: $\{Q\}$. Generate $K$ sub-questions that might help you confirm whether the answer to the target question is ’$\{a\}$’.
@@
[…]
$\mathcal{E}_{RR}$: A list of reliable and relevant evidences [$e_{1},e_{2},...$]
Premise: {evidence statements in $\{\text{evidence statements in }\mathcal{E}_{RR}\text{ concatenated together}\}$
@@ Appendix C OK-VQA and Sherlock Task Details
Is the given inference true for the given image or not? { Inference } Options: yes, no